OpenStudy (anonymous):

In the quadratic equation x2 - 8x + a = 0; what is the value, a, in order to form a perfect square?

OpenStudy (tkhunny):

Have you considered \((a+b)^{2} = a^{2} + 2ab + b^{2}\) ?

OpenStudy (anonymous):

can you explain the problem plz cause i dont really get it @tkhunny

OpenStudy (tkhunny):

That is the action of a perfect square. That's how they work. You have the first two pieces. How do you find the third?

OpenStudy (tkhunny):

Explore it a little: \((x-1)^{2} = x^{2} - 2x + 1\) \((x-2)^{2} = x^{2} - 4x + 4\) \((x-3)^{2} = x^{2} - 6x + 9\) Are we getting close? \((x-4)^{2} =\) ??

OpenStudy (anonymous):

im still confuse to be honest im sorry